OpenStudy (anonymous):

a rectangle has a diagonal that measures 10 centimeters and a length that is 2 centimeters longer than the width. What is the width of the rectangle in centimeters?

OpenStudy (anonymous):

6

OpenStudy (anonymous):

 you could use Pythagorean theorem

OpenStudy (anonymous):

\[\sqrt{x^2+y^2}=10\] \[x=y+2\] \[2y^2+4y+4=100\] \[y^2+2y-48=0\]

OpenStudy (anonymous):

\[10^2-(2+w)^2=w^2\] \[100-(4+4w+w^2)=w^2\] \[100-4-4w-w^2=w^2\] \[2w^2+4w-96=0\] \[w^2+2w-48=0\] \[(w+8)(w-6)\] \[w=-8, w=6\] Reject -8 as it is negative and you cannot have a negative length.